3. Gains from trade Consider two neighboring island countries called Dolorium and Contente. They each have 4 milion labor hours available per month that they can use to produce jeans, corn, or a combination of both. The following table shows the amount of jeans or corn that can be produced using 1 hour of labor. Jeans Corn (Bushels per hour of labor) Country Dolorium Contente (Pairs per hour of labor) 8 Initially, suppose Contente uses 1 million hours of labor per month to produce jeans and 3 million hours per month to produce corn, while Dolorium uses 3 million hours of labor per month to produce jeans and 1 million hours per month to produce corn. Consequently, Dolorium produces 15 million pairs of jeans and 20 million bushels of corn, and Contente produces 8 million pairs of jeans and 48 million bushels of corn. Assume there are no other countries willing to trade goods, so, in the absence of trade between these two countries, each country consumes the amount of jeans and corn it produces Doloriums opportunity cost of producing 1 pair of jeans is of corn, and Contentes opportunity cost of producing 1 pair of jeans is has a comparative advantage in the production of jeans, and of corn. Therefore, advantage in the production of corn. has a comparative Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces jeans will produce million pairs per month, and the country that produces corn will produce million bushels per month1/2 bushel 1/4 bushel 2 bushels 4 bushelsDolorium ContenteIn the following table, enter each countrys production decision on the third row of the table (marked Production) Suppose the country that produces jeans trades 18 million pairs of jeans to the other country in exchange for 54 million bushels of corn In the following table, select the amount of each good that each country exports and imports in the boxes across the row marked Trade Action, and enter each countrys final consumption of each good on the line marked Consumption. When the two countries did not specialize, the total production of jeans was 23 million pairs per month, and the total production of corn was 68 million bushels per month. Because of specialization, the total production of jeans has increased by of corn has increased by million pairs per month, and the total production million bushels per month Because the two countries produce more jeans and more corn under specialization, each country is able to gain from trade Calculate the gains from trade-that is, the amount by which each country has increased its consumption of each good relative to the first row of the table. In the following table, enter this difference in the boxes across the last row (marked Increase in Consumption) Dolorium Contente eans Corn eans Corn (Millions of pairs) (Millions of bushels) (Millions of pairs) (Millions of bushels) Without Trade Production 15 20 8 48 Consumption 15 20 8 48 With Trade Production Trade action Consumption Gains from Trade Increase in ConsumptionExports 18 Imports 18Exports 54 Imports 54
Answer #1

Dolorium's opportunity cost of producing 1 pair of jeans is 20/5 = 4 bushel of corn and Contente's opportunity cost of producing 1 pair of jeans is 16/8 =2 bushel of corn. Because Contente has a lower opportunity cost for the production of jeans. Therefore, Contente has a comparative advantage in the production of jeans and Dolorium has a comparative advantage in the production of corn

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage , producing only that good . In this case,the country that produces jeans i.e Contente will produce (4)(8) = 32 million pairs per month And the country that produces corn i.e Dolorium will produce 4(20)=80 million bushels per month.

Suppose the country that produces corn trades 18 million pairs of jeans  to the other country in exchange for 54 million bushels of corn.

DOLORIUM CONTENTE
Jeans (milions of Pairs) Corn (Millions of bushels) Jeans (Millions of pairs) corn (Millions of bushels)
Without trade Production 15 20 8 48
Consumption 15 20 8 48
With Trade Production 0 80 32 0
Trade action Import 18 Export 54 Export 18 Import 54
Consumption 18 (80-54)= 26

(32-18)=14

54
Gains from trade Increase in consumption (18-15)=3 (26-22)=4 (14-8)=6 (54-48)=

When the two countries did not specialize ,the total production of jeans was 23 million pairs per month and the total production of corn was 68 millions bushels per month. Because of specialization ,the total production of jeans was 32 million pairs per month and total production of corn was 80 million bushels per month. This implies that because of specialization , total production of jeans  has increased by (32-23)=9 million pairs per month and total production of corn has increased by (80-68)=12 million bushels per month.